OpenStudy (amit69):

Find the effective annual interest rate r of the given annual interest rate. Round your answer to the nearest 0.01%. 10% compounded monthly. r= ?

OpenStudy (2500161):

0.15 / 365 = k (15 + r)^(15/365) - 15 = r

OpenStudy (amit69):

@2500161 Thank you very much ^_^

OpenStudy (2500161):

Your welcome @amit69